# Repoformer: Selective Retrieval for Repository-Level Code Completion

### v1

@@ Discussion
Building on our study, future research may consider further leveraging Repoformer for speeding up large LMs in settings such as speculative decoding (Chen et al., 2023). Besides, the current function completion performance could be further improved by adapting selective RAG to dynamically invoking multiple retrievals during the generation. Moreover, tailoring retrieval policies to individual repositories by leveraging their unique characteristics could improve the effectiveness of our selective retrieval framework, aligning with our goal of developing smarter, more efficient RAG-based code completion solutions.
@@
Our research introduces a novel approach to repository-level code completion that significantly enhances efficiency and accuracy by employing selective retrieval, reducing unnecessary computational waste and contributing to more sustainable software development practices. While promising in streamlining development workflows and potentially applicable in various domains, it is important to consider the implications of increased automation in software development, programming education, and the potential for inadvertent biases. Ensuring the ethical use and ongoing evaluation of such code automation technologies is crucial to maximizing their societal benefits while mitigating risks. In this work, we mainly rely on open-sourced, permissively-licensed repositories (the Stack, CrossCodeEval) and models (StarCoder, CodeGen) to perform the experiments. However, as mentioned by Ding et al. (2023), some of the repositories of RepoEval are with non-permissive licenses. We rely on the dataset and code distributed by the original RepoEval authors to perform the experiment and do not re-distribute the dataset or adapt it for other purposes.
@@ Appendix B Why infilling?
To illustrate, we investigate the effect of directly providing $X_{r}$ in the prompt for CodeGen-Mono 16B and StarCoder on current-file code completion and retrieval-augmented code completion. Figure 6 presents the performance on RepoEval with different types of contexts provided in the prompt. Whether cross-file contexts are present or not, providing right contexts can greatly improve the code completion performance. The gain is consistent for both API and function completion. Compared to CodeGen, StarCoder can better leverage the right context to generate more accurate code. Overall, we observe that leveraging the entire right context to perform infilling represents a much stronger baseline. Therefore, in this paper we have exclusively focused on the infilling setting with StarCoder.
@@ C.1 Trial Retrieval
Figure 7 presents the selective RAG performance of StarCoder under different budgets.We observe that the retrievers’ similarity scores serve as a promising signal for deciding whether the retrieved information can improve the RAG performance. For most retrievers and tasks, the performance of full retrieval could be reached with at most 60% retrieval budget. This trend also aligns with the remark in Zhang et al. (2023) on the correlation between in-repository duplication and the gain from $CC$. However, it is worth noting that this strategy brings no latency gain as it still implements invariable retrieval. In addition, the knowledge of whether the LM could be benefited by the retrieved context is not leveraged.
@@ C.2 Trial Generation
  • entropy: the sequence-level entropy as used in Li et al. (2023a). We estimate the entropy by performing vanilla sampling for 20 times.
@@
Appendix D Data Creation for Repoformer Training and CCEval
@@
        $Y\leftarrow$ Cut a sub-chunk from $s$ that spans $k$ consecutive lines
@@
           $\mathcal{Q}\leftarrow$ Concatenate(last $5k$ lines of $X_{l}$, $Y$, first $5k$ lines of $X_{r}$)
@@
For the 240k chunk completion and 120k function completion instances, we plot the performance change after providing $CC$ in Figure 9. In total, 30.18% chunk completion instances and 35.16% function completion instances are labeled with positive. The average length of $Y$ is 3.53 lines for chunk completion and 11.77 lines for function completion.
@@
CCEval Benchmarking Dataset Creation
@@
  • Data analysis In Table 6, we present the basic statistics of RepoEval and CCEval.
RepoEval CCEval
@@
Table 6: Descriptive statistics of RepoEval and CCEval. For $|Y|$, $|X_{l}|$, and $|X_{r}|$, we report both the number of lines as well as the number of tokens (using the StarCoder tokenizer) in the groundtruth, left context, and the right context.
@@ E.1 Calibration of Repoformer’s Selective Retrieval Prediction
E.2 Results on CrossCodeEval: Repoformer-7B and Multilingual Repoformer
@@
Repoformer-7B
We apply the Repoformer training scheme on Repoformer-7B with the dataset created for Repoformer-1B/3B. We keep the training infrastructure and hyperparameters the same.
@@
  • Repoformer-7B. We observe the same trends as we have observed on the 1B and 3B model. Selective retrieval with Repoformer is able to outperform invariable retrieval with the original StarCoder model.
@@
Repoformer-1B $P$(<cc>) 71.29 62.81 75.12 67.16 83.08 74.24 69.90 64.07
Repoformer-3B $P$(<cc>) 74.57 66.86 78.40 71.26 85.92 78.62 73.70 68.66
Repoformer-7B $P$(<cc>) 75.34 68.27 78.90 72.35 83.80 76.88 73.59 69.10
Repoformer-M-1B $P$(<cc>) 71.55 62.89 75.92 67.86 84.44 76.00 70.07 64.41
Repoformer-M-3B $P$(<cc>) 73.80 66.72 77.68 71.01 85.31 77.70 72.51 67.06
Repoformer-M-7B $P$(<cc>) 75.35 67.88 79.11 72.82 86.53 79.77 74.60 70.01
Table 7: Evaluation results on CrossCodeEval. We report edit similar for code matching as well as the F1 score for identifier matching.
@@ E.3 Repoformer’s Robustness to the Retriever Choice
In this section, we investigate the performance of Repoformer with the cosine similarity of UniXcoder embedding (Guo et al., 2022) as the retriever instead of Jaccard similarity. As shown in Table 8, we are able to observe similar patterns compared to Table 3: selective retrieval is able to improve both the accuracy and the latency of the entire RAG system. In addition, as retrieval consumes a larger proportion of latency than when sparse retriever is used, selective retrieval brings more substantial performance gains, with $P$(<cc>) bringing more than 70% speedup.
Selective Policy API Completion Line Completion
@@
invariable retrieval 71.69 $100\%$ $0\%$ 75.25 $100\%$ 0%
self-selection 70.82 $18\%$ $71\%$ 73.70 $19\%$ $71\%$
1B $P$(<cc>) 72.39 $61\%$ $33\%$ 75.65 $62\%$ $33\%$
invariable retrieval 74.48 $100\%$ $0\%$ 78.24 $100\%$ 0%
self-selection 73.26 $19\%$ $65\%$ 76.74 $20\%$ $66\%$
3B $P$(<cc>) 74.69 $78\%$ $21\%$ 78.63 $74\%$ $31\%$
Table 8: RAG performance of Repoformer with two self-selective RAG paradigms and dense retrieval used instead of Jaccard similarity. %RAG = ratio of instances where RAG is performed. SU = Speedup compared to invariable retrieval. Compared to the invariable retrieval baseline, the $P$(<cc>) strategy consistently demonstrates gains in both accuracy and latency. The self-selection strategy shows much larger latency gains with a small performance degradation. Compared to sparse retrieval, we observe more substantial latency gains.

### v2

@@ Discussion
Building upon Repoformer, future research may consider several important directions:
  1. 1.
    Further speeding up large LMs. Beyond as a selective retrieval policy, Repoformer has the potential to serve as an effective plug-in draft model in settings such as speculative decoding (Chen et al., 2023).
  2. 2.
    More effective function completion. To enable a good scalability, we used lexical similarity as the signal for training label creation. Although this heuristics enables improvements in function completion evaluation, designing a more accurate and scalable labeling approach is an important future direction.
  3. 3.
    Personalized retrieval. We apply a uniform selective policy across repositories. However, certain repositories could be inherently more RAG-friendly by exhibiting a higher level of duplication (Zhang et al., 2023). Adapting the selective RAG paradigm towards accurate personalized policies is an important direction.
Acknowledgement
We express our gratitude to anonymous reviewers for their valuable suggestions to improve the quality of the paper. The authors also thank Amita Kamath and Po-Nien Kung for their constructive feedback provided during the paper writing process. Additionally, we would like to express gratitude to some other team members from Amazon CodeWhisperer and UCLANLP for their insightful discussions, which have contributed to the refinement of our work.
@@ Acknowledgement
Our research introduces a novel approach to repository-level code completion that significantly enhances efficiency and accuracy by employing selective retrieval, reducing unnecessary computational waste, and contributing to more sustainable software development practices. Although promising in streamlining development workflows and potentially applicable in various domains, it is important to consider the implications of increased automation in software development, programming education, and the potential for inadvertent biases. Ensuring the ethical use and ongoing evaluation of such code automation technologies is crucial to maximize their societal benefits while mitigating risks. In addition, as a general infrastructure, it is important to design additional mechanisms that prevent RAG systems from revealing sensitive data in the retrieval database.
In this work, we mainly rely on open-sourced, permissively-licensed repositories (the Stack, CrossCodeEval) and models (StarCoder, CodeGen) to perform the experiments. However, as mentioned by Ding et al. (2023), some of the repositories of RepoEval are with non-permissive licenses. We rely on the dataset and code distributed by the original RepoEval authors to perform the experiment and do not re-distribute the dataset or adapt it for other purposes.
@@ Appendix B Why infilling?
To illustrate, we investigate the effect of directly providing $X_{r}$ in the prompt for CodeGen-Mono 16B and StarCoder on current-file code completion and retrieval-augmented code completion. Figure 7 presents the performance on RepoEval with different types of contexts provided in the prompt. Whether cross-file contexts are present or not, providing right contexts can greatly improve the code completion performance. The gain is consistent for both API and function completion. Compared to CodeGen, StarCoder can better leverage the right context to generate more accurate code. Overall, we observe that leveraging the entire right context to perform infilling represents a much stronger baseline. Therefore, in this paper we have exclusively focused on the infilling setting with the StarCoder family.
@@ C.1 Trial Retrieval
Figure 8 presents the selective RAG performance of StarCoder under different budgets.We observe that the retrievers’ similarity scores serve as a promising signal for deciding whether the retrieved information can improve the RAG performance. For most retrievers and tasks, the performance of full retrieval could be reached with at most 60% retrieval budget. This trend also aligns with the remark in Zhang et al. (2023) on the correlation between in-repository duplication and the gain from $CC$. However, it is worth noting that this strategy brings no latency gain as it still implements always retrieving. In addition, the knowledge of whether the LM could be benefited by the retrieved context is not leveraged.
@@ C.2 Trial Generation
  • entropy: the sequence-level entropy as used in Li et al. (2023a). We estimate the entropy by performing vanilla sampling for 20 times without any temperature scaling or distribution truncation.
@@
Appendix D Data Creation for Repoformer Training and CrossCodeLongEval
@@
        $Y\leftarrow$ Randomly cut a sub-chunk from $s$ that spans $k$ consecutive lines
@@
           $\mathcal{Q}\leftarrow$ Concatenate(last $5k$ lines of $X_{l}$, $Y$, first $5k$ lines of $X_{r}$) // query formation
@@
For the 240k chunk completion and 120k function completion instances, we plot the performance change after providing $CC$ in Figure 10. In total, 30.18% chunk completion instances and 35.16% function completion instances are labeled with positive (i.e., retrieval should be triggered). The average length of $Y$ is 3.53 lines for chunk completion and 11.77 lines for function completion.
@@
CrossCodeLongEval Construction
@@
  • Data analysis In Table 6, we present the basic statistics of RepoEval and CrossCodeLongEval.
RepoEval CrossCodeLongEval
@@
Table 6: Descriptive statistics of RepoEval and CrossCodeLongEval. For $|Y|$, $|X_{l}|$, and $|X_{r}|$, we report both the number of lines as well as the number of tokens (using the StarCoder tokenizer) in the groundtruth, left context, and the right context.
@@ E.1 Calibration of Repoformer’s Selective Retrieval Prediction
E.2 CrossCodeEval and Multilingual Repoformer
@@
[…]
Repoformer-M-7B SelectiveT 75.35 67.88 79.11 72.82 86.53 79.77 74.60 70.01
Table 7: Evaluation results on CrossCodeEval. We report edit similarity for code matching as well as the F1 score for identifier matching. The best scores across all models are boldfaced.
@@ E.3 Repoformer’s Robustness to the Retriever Choice
In this section, we investigate the performance of Repoformer with the cosine similarity of UniXcoder embedding (Guo et al., 2022) as the retriever instead of Jaccard similarity. As shown in Table 8, we are able to observe similar patterns compared to Table 3: selective retrieval is able to improve both the accuracy and the latency of the entire RAG system. In addition, as retrieval consumes a larger proportion of latency than when sparse retriever is used, selective retrieval brings more substantial performance gains, with threshold selection bringing more than 70% speedup.
Model RAG Policy API Completion Line Completion
@@
Always 71.69 $100\%$ $0\%$ 75.25 $100\%$ 0%
SelectiveG 70.82 $18\%$ $71\%$ 73.70 $19\%$ $71\%$
Repoformer-1B SelectiveT 72.39 $61\%$ $33\%$ 75.65 $62\%$ $33\%$
Always 74.48 $100\%$ $0\%$ 78.24 $100\%$ 0%
SelectiveG 73.26 $19\%$ $65\%$ 76.74 $20\%$ $66\%$
Repoformer-3B SelectiveT 74.69 $78\%$ $21\%$ 78.63 $74\%$ $31\%$
Table 8: RAG performance of Repoformer with two self-selective RAG paradigms and dense retrieval used instead of Jaccard similarity. %RAG = ratio of instances where RAG is performed. SU = Speedup compared to always retrieving. Compared to the always retrieving baseline, the SelectiveT strategy consistently demonstrates gains in both accuracy and latency. The SelectiveG strategy shows much larger latency gains with a small performance degradation. Compared to sparse retrieval, we observe more substantial latency gains.